OpenStudy (anonymous):

The floor of a rectangular room will be tiled with square tiles that measure 3 inches on each side. The room is 109 inches long and 103 inches wide. How many tiles, including parts of tiles, are needed to complete the job? A. 1,248 tiles B. 1,247 tiles C. 1,224 tiles D. 1,295 tiles

OpenStudy (anonymous):

E

OpenStudy (anonymous):

theres not an E

OpenStudy (anonymous):

What are you having trouble with? Can you show your work?

OpenStudy (anonymous):

idk how to work it out

OpenStudy (anonymous):

Do you know the area of the floor?

OpenStudy (anonymous):

That's how much space the floor has.

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

Ok the area of the floor. which simply is the area of a rectangle is the width times the height. Or in this case 109 * 103 What's 109* 103?

OpenStudy (anonymous):

11,227

OpenStudy (anonymous):

S rectangular = 103*109 =11227 S tiled =3^2=9 11227 : 9 =1247

OpenStudy (anonymous):

@Palmo4ka don't snipe. No one really cares if you know how to solve the problem.

OpenStudy (anonymous):

thats rude Romero

OpenStudy (anonymous):

@Palmo4ka..thanks